OpenStudy (anonymous):

[7.04] Choose the correct simplification of 3 over x to the power of negative 7. Select one: a. x to the 7th power over 3 b. 3 over x to the 7th power c. 3x7 d. Already simplified.

OpenStudy (johnweldon1993):

\[\huge (\frac{3}{x^{-7}}) = 3 \times x^7 = \space ?\]

OpenStudy (anonymous):

c?

OpenStudy (johnweldon1993):

Mmhmm :)

OpenStudy (anonymous):

Yay!